# EXAONE 4.0: Unified Large Language Models Integrating Non-reasoning and Reasoning Modes

### v1

@@ D.1 HELMET
We include the HELMET benchmark [66] in our evaluation to systematically assess models’ long-context capabilities across both synthetic and real-world tasks. HELMET is designed as a comprehensive suite of diverse, application-centric tasks and addresses key limitations of prior benchmarks, such as inadequate input lengths, over-reliance on retrieval-style setups, and unreliable evaluation metrics. Crucially, it covers a wide spectrum of long-context challenges, including information recall, multi-hop retrieval, in-context generalization, and long-input generation, making it a well-suited benchmark for evaluating models’ ability to process and reason over extended sequences in practical settings.
@@ D.2 RULER
We evaluate our model’s long-context capabilities using the RULER benchmark [22], a synthetic evaluation suite designed to assess various aspects of long-context understanding beyond simple retrieval. RULER consists of diverse task categories, including retrieval, multi-hop tracing, aggregation, and question answering, and supports flexible configurations for context length and task complexity. The performance of our models across different sequence lengths is summarized in Table 10.
@@ D.3 LongBench
LongBench [5] has been suggested as a bilingual benchmark to assess long context comprehension in English and Chinese. We focus on the English subsets, specifically Single-doc QA, Multi-doc QA, Summarization, and Few-shot Learning.
The Single-doc QA task covers datasets such as NarrativeQA [27], Qasper [7], and MultiFieldQA-EN [5]. For the Multi-doc QA task, we employ benchmarks including HotpotQA [64], 2WikiMultihopQA [20], and MuSiQue [57]. The Summarization task utilizes datasets like GovReport [23], QMSum [68], and MultiNews [11]. For the Few-shot Learning task, we draw from TREC [33] and TriviaQA [25]. All evaluations follow the official protocols and metrics defined in LongBench.

### v2

@@ D.1 HELMET
We include the HELMET benchmark [yen2025helmetevaluatelongcontextlanguage] in our evaluation to systematically assess models’ long-context capabilities across both synthetic and real-world tasks. HELMET is designed as a comprehensive suite of diverse, application-centric tasks and addresses key limitations of prior benchmarks, such as inadequate input lengths, over-reliance on retrieval-style setups, and unreliable evaluation metrics. Crucially, it covers a wide spectrum of long-context challenges, including information recall, multi-hop retrieval, in-context generalization, and long-input generation, making it a well-suited benchmark for evaluating models’ ability to process and reason over extended sequences in practical settings.
@@ D.2 RULER
We evaluate our model’s long-context capabilities using the RULER benchmark [hsieh2024rulerwhatsrealcontext], a synthetic evaluation suite designed to assess various aspects of long-context understanding beyond simple retrieval. RULER consists of diverse task categories, including retrieval, multi-hop tracing, aggregation, and question answering, and supports flexible configurations for context length and task complexity. The performance of our models across different sequence lengths is summarized in Table 10.
@@ D.3 LongBench
LongBench [bai2024longbenchbilingualmultitaskbenchmark] has been suggested as a bilingual benchmark to assess long context comprehension in English and Chinese. We focus on the English subsets, specifically Single-doc QA, Multi-doc QA, Summarization, and Few-shot Learning.
The Single-doc QA task covers datasets such as NarrativeQA [kovcisky2018narrativeqa], Qasper [dasigi-etal-2021-dataset], and MultiFieldQA-EN [bai2024longbenchbilingualmultitaskbenchmark]. For the Multi-doc QA task, we employ benchmarks including HotpotQA [yang-etal-2018-hotpotqa], 2WikiMultihopQA [ho-etal-2020-constructing], and MuSiQue [trivedi-etal-2022-musique]. The Summarization task utilizes datasets like GovReport [huang-etal-2021-efficient], QMSum [zhong-etal-2021-qmsum], and MultiNews [fabbri-etal-2019-multi]. For the Few-shot Learning task, we draw from TREC [li-roth-2002-learning] and TriviaQA [joshi-etal-2017-triviaqa]. All evaluations follow the official protocols and metrics defined in LongBench.